# Structure-Informed Estimation for Pilot-Limited MIMO Channels via Tensor Decomposition

### v1

@@ 3.3 Tucker Decomposition
While CP enforces strict rank-one separability across all modes, a more flexible multilinear representation is provided by the Tucker decomposition [24],
@@
where $\mathcal{G}\in\mathbb{C}^{R_{r}\times R_{t}\times R_{f}}$ is a low-dimensional core tensor, and $\mathbf{U}_{r}\in\mathbb{C}^{N_{r}\times R_{r}}$, $\mathbf{U}_{t}\in\mathbb{C}^{N_{t}\times R_{t}}$, and $\mathbf{U}_{f}\in\mathbb{C}^{N_{f}\times R_{f}}$ are factor matrices defining low-dimensional subspaces along each mode.
Tucker generalizes CP, which is recovered when the core tensor $\mathcal{G}$ is superdiagonal with $R_{r}=R_{t}=R_{f}$, by relaxing strict separability across modes. The core tensor captures inter-mode interactions that cannot be expressed as independent rank-one components, enabling representation of richer spatial–spectral correlations. From a physical perspective, this flexibility allows Tucker models to capture clustered multipath and moderately diffuse scattering, where multiple rays share correlated angular and delay characteristics. The multilinear ranks $(R_{r},R_{t},R_{f})$ thus control a trade-off between compression and fidelity: smaller ranks yield stronger dimensionality reduction, while larger ranks enable finer modeling of scattering richness and model mismatch.
@@ 3.4 Algebraic Priors from Propagation Physics
  • Structured factor matrices, capturing array-induced spatial correlations (e.g., approximately Vandermonde structure in uniform linear arrays);
@@
The low-rank property has been extensively exploited through matrix completion and matrix sensing formulations for mmWave and terahertz (THz) systems [25], complementing the tensor-based approach adopted in this work. Unlike geometric priors that rely on site-specific environmental descriptions or ray-tracing information, these algebraic constraints arise intrinsically from propagation physics and array structure, and therefore persist across diverse deployment scenarios. As a result, tensor representations provide robust, interpretable, and physically grounded priors for pilot-limited channel estimation, consistent with prior tensor-based formulations in wireless communications [17].
@@ 4 Structure-Informed Learning Framework
This section presents the proposed structure-informed learning framework for pilot-limited MIMO channel estimation. The framework is designed to explicitly combine algebraic priors induced by propagation physics with data-driven residual learning in a unified estimation pipeline. A low-rank tensor completion stage first enforces intrinsic multilinear structure arising from sparse multipath propagation, array geometry, and frequency coherence, yielding a physically consistent coarse channel estimate. A lightweight neural network is then employed to learn residual components that are not captured by the idealized low-rank model, such as diffuse scattering effects, hardware non-idealities, and model mismatch. By decoupling structure enforcement from residual learning, the proposed approach achieves improved robustness and generalization in realistic pilot-limited regimes while retaining interpretability and computational efficiency. Fig. 1 illustrates the proposed two-stage estimation pipeline.
@@ 4.1 Problem Formulation
the goal is to estimate the unknown channel tensor $\mathcal{H}$ from incomplete and noisy measurements. This is posed as
$\widehat{\mathcal{H}}=\arg\min_{\mathcal{H}}\left\|\mathcal{P}_{\Omega}(\mathcal{H})-\mathcal{Y}_{\Omega}\right\|_{F}^{2},$ (12)
where $\mathcal{P}_{\Omega}(\cdot)$ denotes projection onto the observed index set $\Omega$ and $\mathcal{N}$ is i.i.d. complex Gaussian noise.
@@ 4.2 Tensor Completion via Alternating Projection
Low-rank structure is enforced via an alternating projection scheme that iterates between: (i) a low-rank projection, computed using truncated HOSVD for Tucker models [24]; and (ii) a data-consistency projection, which enforces agreement with the observed pilot samples. For the baseline labeled “CP”, the default implementation does not run masked CP alternating least squares (CP-ALS); instead, it uses a stable Tucker-completion step under partial observations and returns the completed tensor. A weighted CP-ALS formulation is implemented but not used by default due to numerical instability in sparse, noisy settings. The resulting estimate satisfies both the algebraic low-rank constraint and consistency with the observed data.
@@
While low-rank tensor models capture the dominant physical structure induced by sparse multipath propagation, they inevitably rely on idealized assumptions that may be violated in practice. In realistic systems, effects such as diffuse scattering, clustered multipath, hardware impairments, and calibration errors introduce structured deviations that cannot be fully represented by CP or Tucker models alone. To address this limitation, the proposed framework augments algebraic tensor completion with a lightweight neural residual learning stage. Importantly, the neural network does not merely perform noise suppression; instead, it learns systematic residual components arising from model mismatch, complementing the physics-inspired tensor model rather than replacing it. This hybrid design places the proposed method within the class of model-aided or physics-informed learning approaches, where algebraic structure provides the primary inductive bias and data-driven learning is restricted to correcting non-idealities beyond the expressive power of the underlying mathematical model.
Algorithm 1 details the Tucker completion procedure. Let $\mathcal{Y}$ denote the observed tensor (with unobserved entries set to zero) and let $\mathcal{M}\in\{0,1\}^{N_{r}\times N_{t}\times N_{f}}$ be the binary mask indicating observed entries ($\mathcal{M}]_{ijk}=1$ if $(i,j,k)\in\Omega$). The algorithm seeks a tensor $\widehat{\mathcal{H}}$ of multilinear rank $(R_{r},R_{t},R_{f})$ consistent with the observations.
@@
$\mathcal{X}^{(0)}\leftarrow\mathcal{Y}$ {Zero-fill unobserved entries}
@@
Stopping criteria
Iteration terminates when either: (i) the relative change in observed entries falls below tolerance $\epsilon$, or (ii) the maximum number of iterations $T$ is reached. In all experiments, $\epsilon=10^{-6}$ and $T=20$ unless otherwise specified.
Convergence properties
Alternating projection for Tucker completion is a block coordinate descent method on a non-convex feasibility problem and typically converges to a stationary point under mild conditions [26]. Global optimality is not guaranteed due to non-convexity of the low-rank constraint. For CP-ALS, the landscape is similarly non-convex, and the algorithm may converge to local minima. Multiple random initializations refer to different initial factor matrices generated deterministically from a fixed base seed.
Rank selection
The multilinear ranks $(R_{r},R_{t},R_{f})$ for Tucker and the CP rank $R$ are assumed known or estimated a priori. In practice, rank selection can be performed via model-order criteria such as minimum description length (MDL) or Akaike information criterion (AIC) applied to singular values of tensor unfoldings [23]. For synthetic experiments, ranks are set to match the number of propagation paths ($L=5$). For DeepMIMO channels, where effective rank is empirical and environment-dependent, ranks are chosen conservatively based on singular value analysis of representative channel realizations. When assumed rank underestimates true channel complexity, the residual $\Delta\mathcal{H}=\mathcal{H}-\mathcal{H}_{\text{LR}}$ increases; the neural residual component is designed to compensate for such model mismatch, as demonstrated in Experiment 5.4.
@@ 4.3 Computational Complexity
[…]
@@ 4.4 Hybrid Tensor–Neural Extension
Although tensor decompositions capture dominant low-dimensional structure induced by sparse multipath propagation, residual errors arise from diffuse scattering, hardware impairments, or model mismatch. To account for these effects, the channel is modeled as
@@
where $\mathcal{T}(\bm{\theta})$ denotes a low-rank tensor component and $\Delta\mathcal{H}_{\text{NN}}(\bm{\phi})$ a lightweight neural residual. Learning is thus restricted to a correction term rather than the full channel, reducing sample complexity, mitigating overfitting, and improving generalization.
Estimation proceeds in two stages. First, the structured component $\mathcal{T}(\bm{\theta})$ is obtained by solving the rank-constrained problem
$\min_{\mathcal{T}}\left\|\mathcal{Y}_{\Omega}-\mathcal{P}_{\Omega}(\mathcal{T})\right\|_{F}^{2}\quad\text{subject to fixed Tucker or CP rank},$ (14)
via truncated HOSVD with alternating projections for Tucker. Second, the neural network learns the residual by minimizing the supervised mean-squared error (MSE) between predicted and true residuals,
@@
where $\mathcal{H}_{\text{LR}}$ is the low-rank reconstruction and $\mathcal{H}$ the ground-truth channel. This two-stage approach aligns with model-driven and algorithm-unrolling paradigms in signal processing [27, 28, 29].
@@ 4.4.1 Neural Architecture and Training
The residual term $\Delta\mathcal{H}_{\text{NN}}$ is modeled using a lightweight three-dimensional U-Net architecture [30], denoted Residual3DUNet, specifically designed for residual prediction. The network operates directly on the spatial–frequency channel tensor and exploits encoder–decoder symmetry with skip connections to combine global context and local detail while maintaining a limited parameter count. Table 1 summarizes the layer-by-layer architecture, including channel dimensions and convolutional parameters.
@@
The input consists of the observed tensor $\mathcal{Y}_{\Omega}$ represented by separate real and imaginary channels, yielding $C_{\text{in}}=2$ input channels, optionally augmented with a binary pilot mask ($C_{\text{in}}=3$). A single encoder–decoder level is employed. The base channel width is denoted by $C$, with $C=16$ in all experiments unless otherwise stated.
The encoder applies two $3\times 3\times 3$ convolutional layers with padding 1, each followed by ReLU activation, and performs downsampling via a $2\times 2\times 2$ strided convolution doubling the channel dimension. The second encoder block begins with ReLU, followed by a $3\times 3\times 3$ convolution and another ReLU. The decoder mirrors this structure using transposed convolution for upsampling, concatenation with corresponding encoder features through a skip connection, and two $3\times 3\times 3$ convolutions with ReLU activations. The final layer projects features back to the input channel dimension, producing the residual estimate $\widehat{\mathcal{R}}$. When necessary, trilinear interpolation aligns tensor dimensions prior to skip-connection concatenation. No batch normalization, dropout, or other explicit regularization layers are employed.
The reconstructed channel is obtained as
@@
Training uses the Adam optimizer [31] with fixed learning rate $10^{-3}$ and batch size 2 (unless otherwise stated). No learning-rate decay or warm-up is used. Early stopping is applied only to the ResNet baseline and is disabled by default; the hybrid Tensor–NN is trained for a fixed number of epochs.
@@
All neural networks are trained using the Adam optimizer with fixed learning rate $10^{-3}$ and no learning-rate decay or warm-up. Network weights are initialized using PyTorch default initialization schemes, corresponding to Kaiming/He uniform initialization for convolutional and transposed convolutional layers, with biases initialized to zero when present [32].
No explicit data augmentation is applied. Training variability is provided by random channel generation and, when enabled, by sampling pilot ratios and SNR values over predefined ranges (multi-pilot and multi-SNR training). For the proposed hybrid Tensor–NN model, training is performed for a fixed number of epochs without early stopping.
@@ 4.4.2 Residual Structure and Architectural Rationale
The U-Net architecture is well matched to this residual structure: its encoder captures coarse global context through progressive downsampling, while skip connections preserve fine-scale localized information. This design enables efficient reconstruction of multi-scale residual patterns with a small parameter budget, a property proven effective in inverse problems and denoising tasks involving structured residuals.
@@ 5 Simulation Results
This section evaluates pilot-limited wideband MIMO channel estimation under synthetic and ray-tracing-based propagation models. The goal is to quantify how intrinsic tensor structure—in particular, low multilinear rank induced by sparse scattering—enables reliable reconstruction of the wideband channel tensor $\mathcal{H}$ from noisy and incomplete pilot observations [17].

### v2

@@ 3.2 Tucker Decomposition
While CP enforces strict rank-one separability across all modes, the Tucker decomposition [24] provides a more flexible multilinear representation,
@@
where $\mathcal{G}\in\mathbb{C}^{R_{r}\times R_{t}\times R_{f}}$ is a low-dimensional core tensor capturing inter-mode coupling. The integers $(R_{r},R_{t},R_{f})$ denote the multilinear ranks of $\mathcal{H}$ along the receive, transmit, and frequency modes, respectively. The factor matrices $\mathbf{U}_{r}\in\mathbb{C}^{N_{r}\times R_{r}}$, $\mathbf{U}_{t}\in\mathbb{C}^{N_{t}\times R_{t}}$, and $\mathbf{U}_{f}\in\mathbb{C}^{N_{f}\times R_{f}}$ contain orthonormal columns spanning the dominant subspaces associated with each mode.
Tucker generalizes CP, which the formulation recovers when the core tensor $\mathcal{G}$ is superdiagonal with $R_{r}=R_{t}=R_{f}$, by relaxing strict separability across modes. The core tensor captures inter-mode interactions that independent rank-one components cannot express, enabling representation of richer spatial–spectral correlations. From a physical perspective, this flexibility allows Tucker models to capture clustered multipath and moderately diffuse scattering, where multiple rays share correlated angular and delay characteristics. The multilinear ranks $(R_{r},R_{t},R_{f})$ thus control a trade-off between compression and fidelity: smaller ranks yield stronger dimensionality reduction, while larger ranks enable finer modeling of scattering richness and model mismatch.
Fig. 1 illustrates the structural difference between CP and Tucker decompositions.
Figure 1: CP versus Tucker decompositions of the channel tensor $\mathcal{H}\in\mathbb{C}^{N_{r}\times N_{t}\times N_{f}}$. Left: CP expresses $\mathcal{H}$ as a sum of $R$ rank-one components ($\mathbf{u}_{r}\circ\mathbf{v}_{r}\circ\mathbf{w}_{r}$), enforcing strict mode separability. Right: Tucker factorizes $\mathcal{H}$ into a core tensor $\mathcal{G}$ and mode matrices ($\mathcal{G}\times_{1}\mathbf{U}_{r}\times_{2}\mathbf{U}_{t}\times_{3}\mathbf{U}_{f}$), capturing inter-mode correlations.
@@ 3.3 Algebraic Priors from Propagation Physics
  • Structured factor matrices, capturing array-induced spatial correlations;
@@
The literature has extensively exploited the low-rank property through matrix completion and matrix sensing formulations for mmWave and terahertz (THz) systems [25], complementing the tensor-based approach this paper adopts. Unlike geometric priors that rely on site-specific environmental descriptions or ray-tracing information, these algebraic constraints arise intrinsically from propagation physics and array structure, and therefore persist across diverse deployment scenarios. As a result, tensor representations provide robust, interpretable, and physically grounded priors for pilot-limited channel estimation, consistent with prior tensor-based formulations in wireless communications [17].
@@ 4 Structure-Informed Learning Framework
This section presents the proposed structure-informed learning framework for pilot-limited MIMO channel estimation. The framework combines algebraic priors induced by propagation physics with data-driven residual learning in a unified estimation pipeline. A low-rank tensor completion stage first enforces intrinsic multilinear structure arising from sparse multipath propagation, array geometry, and frequency coherence, yielding a physically consistent coarse channel estimate. A lightweight neural network then learns residual components that the idealized low-rank model does not capture, such as diffuse scattering effects, hardware non-idealities, and model mismatch. By decoupling structure enforcement from residual learning, the proposed approach achieves improved robustness and generalization in realistic pilot-limited regimes while retaining interpretability and computational efficiency. Fig. 2 illustrates the proposed two-stage estimation pipeline.
@@ 4.1 Problem Formulation
the goal is to estimate the unknown channel tensor $\mathcal{H}$ from incomplete and noisy measurements. This work poses the problem as
$\begin{split}\widehat{\mathcal{H}}=\;&\arg\min_{\mathcal{H}}\left\|\mathcal{P}_{\Omega}(\mathcal{H})-\mathcal{Y}_{\Omega}\right\|_{F}^{2}\\ &\text{s.t.}\quad\mathrm{rank}_{\mathrm{ml}}(\mathcal{H})=(R_{r},R_{t},R_{f}),\end{split}$ (12)
where $\mathcal{P}_{\Omega}(\cdot)$ denotes projection onto the observed index set $\Omega$, $\mathcal{N}$ is i.i.d. complex Gaussian noise, and $\mathrm{rank}_{\mathrm{ml}}(\mathcal{H})$ denotes the multilinear rank of $\mathcal{H}$.
@@ 4.2 Tensor Completion via Alternating Projection
The alternating projection scheme enforces low-rank structure by iterating between: (i) projection onto a low-rank Tucker manifold via truncated HOSVD [24], and (ii) projection onto the data-consistency set defined by the observed pilot entries. This procedure alternates between structural regularization and measurement agreement.
@@
While low-rank tensor models capture the dominant structure induced by sparse multipath propagation, practical channels exhibit deviations due to diffuse scattering, hardware impairments, and calibration inaccuracies. To account for these effects, the proposed framework augments algebraic tensor completion with a lightweight neural residual stage. The network learns structured residual components associated with model mismatch rather than performing generic denoising. This hybrid formulation follows a model-aided learning perspective in which algebraic structure provides the primary inductive bias and data-driven learning compensates non-idealities.
Algorithm 1 summarizes the Tucker completion procedure. Let $\mathcal{Y}$ denote the observed tensor (zero-filled at unobserved entries) and $\mathcal{M}\in\{0,1\}^{N_{r}\times N_{t}\times N_{f}}$ the sampling mask, where $\mathcal{M}]_{ijk}=1$ if $(i,j,k)\in\Omega$. The objective is to estimate $\widehat{\mathcal{H}}$ with multilinear rank $(R_{r},R_{t},R_{f})$ consistent with the observations. Iterations terminate when $\delta^{(t)}<\epsilon$ or $t=T$; unless otherwise stated, $\epsilon=10^{-6}$ and $T=20$. In all tested configurations, convergence was empirically observed within 10101515 iterations, confirming that $T=20$ provides sufficient margin without introducing unnecessary computation. The tolerance $\epsilon=10^{-6}$ ensures that residual changes on the observed entries are negligible relative to the pilot-set norm, and was verified to be neither too tight nor too loose across the SNR and pilot-density ranges of all experiments.
Alternating projection defines a non-convex feasibility procedure that typically converges to a stationary point under standard regularity conditions [26], without guarantees of global optimality; CP-based formulations exhibit analogous behavior. The framework assumes the multilinear ranks $(R_{r},R_{t},R_{f})$ (Tucker) and the CP rank $R$ known or selected a priori. In practice, rank selection may rely on information criteria (e.g., Akaike Information Criterion (AIC)) or singular-value inspection of tensor unfoldings [23]. Synthetic experiments match ranks to the number of propagation paths ($L=5$), whereas DeepMIMO ranks receive conservative choices based on representative singular-value decay. Underestimated ranks increase the residual $\Delta\mathcal{H}=\mathcal{H}-\mathcal{H}_{\text{LR}}$, motivating the neural residual stage (subsection 5.4).
The Tucker stage performs reconstruction via alternating projections, iterating truncated higher-order singular value decomposition (HOSVD) on the current estimate followed by enforcement of consistency on the observed entries (Algorithm 1). The CP stage performs reconstruction via missing-data-aware weighted alternating least squares (CP-WALS), minimizing a weighted Frobenius norm over the observed entries only. The implementation optionally supports multiple random restarts, selecting the best solution by observed-data fit (NMSE on the pilot set) to reduce initialization sensitivity; when only one restart is used, this selection step is degenerate. Tucker alternating projections maintain a dense iterate and enforce data consistency on the observed set at each iteration, rather than optimizing exclusively on the sparse observation set. In the hybrid Tensor-NN pipeline, the low-rank reconstruction runs first and its output serves as the input prior for neural residual correction.
@@
$\mathcal{X}^{(0)}\leftarrow\mathcal{Y}$
@@ 4.3 Computational Complexity
For CP decomposition with rank $R$, the number of parameters scales as $\mathcal{O}(R(N_{r}+N_{t}+N_{f}))$, compared with $\mathcal{O}(N_{r}N_{t}N_{f})$ for unstructured estimation [23]. For typical dimensions ($N_{r}=N_{t}=32$, $N_{f}=128$, $R=5$), this corresponds to approximately $136\times$ reduction in degrees of freedom. This substantial complexity reduction explains the effectiveness and scalability of structure-informed estimation in pilot-limited MIMO systems, consistent with prior observations [17].
@@ 4.4 Hybrid Tensor–Neural Extension
Although tensor decompositions capture dominant low-dimensional structure induced by sparse multipath propagation, residual errors arise from diffuse scattering, hardware impairments, or model mismatch. To account for these effects, the channel takes the form
@@
where $\mathcal{T}(\bm{\theta})$ denotes a low-rank tensor component and $\Delta\mathcal{H}_{\text{NN}}(\bm{\phi})$ a lightweight neural residual. This decomposition restricts learning to a correction term rather than the full channel, reducing sample complexity, mitigating overfitting, and improving generalization.
Estimation proceeds in two stages. First, the framework obtains the structured component $\mathcal{T}(\bm{\theta})$ by solving the rank-constrained problem (12) via truncated HOSVD with alternating projections for Tucker. Second, the neural network learns the residual by minimizing the supervised mean-squared error (MSE) between predicted and true residuals,
@@
where $\mathcal{H}_{\text{LR}}$ is the low-rank reconstruction and $\mathcal{H}$ the ground-truth channel. This two-stage approach aligns with model-driven and algorithm-unrolling paradigms in signal processing [27, 28, 29]. The two stages are trained sequentially: the tensor completion stage is executed first and its output $\mathcal{H}_{\text{LR}}$ is treated as fixed during neural network training, with no gradient propagation through the tensor stage.
@@ 4.4.1 Neural Architecture and Training
The residual term $\Delta\mathcal{H}_{\text{NN}}$ takes the form of a lightweight three-dimensional U-Net architecture [30], denoted Residual3DUNet. U-Net offers a strong inductive bias in reconstruction tasks, combining multiscale context aggregation with skip connections that preserve fine structural information. This design suits the learning of systematic residual components arising from model mismatch rather than generic denoising. The 3D formulation operates directly on the spatial–frequency tensor, enabling joint modeling of antenna and subcarrier correlations. A lightweight configuration limits parameter count and keeps the tensor model as the primary structural prior. Table 1 and Fig. 3 detail the architecture.
@@
Figure 3: Residual3DUNet architecture ($C_{\text{in}}\in\{2,3\}$, base width $C=16$). Encoder-1 applies two $\text{Conv3D}{+}\text{ReLU}\,(3^{3}/1)$ layers followed by $\text{Conv3D}\,(2^{3}/2)$ downsampling ($C{\to}2C$). Encoder-2 keeps $2C{\to}2C$ at the bottleneck. The decoder upsamples via $\text{ConvTranspose3D}\,(2^{3}/2)$, concatenates encoder skip features, and maps $2C{\to}C$. A final $\text{Conv3D}\,(3^{3}/1)$ projects $C{\to}C_{\text{in}}$, yielding $\Delta\mathcal{H}_{\mathrm{NN}}$.
The input consists of the observed tensor $\mathcal{Y}_{\Omega}$ represented by separate real and imaginary channels, yielding $C_{\text{in}}=2$ input channels, optionally augmented with a binary pilot mask ($C_{\text{in}}=3$). The architecture employs a single encoder–decoder level. The base channel width $C$ equals 16 in all experiments unless otherwise stated.
The encoder applies two $3\times 3\times 3$ convolutional layers with padding 1, each followed by ReLU activation, and performs downsampling via a $2\times 2\times 2$ strided convolution doubling the channel dimension. The second encoder block begins with ReLU, followed by a $3\times 3\times 3$ convolution and another ReLU. The decoder mirrors this structure using transposed convolution for upsampling, concatenation with corresponding encoder features through a skip connection, and two $3\times 3\times 3$ convolutions with ReLU activations. The final layer projects features back to the input channel dimension, producing the residual estimate $\widehat{\mathcal{R}}$. When necessary, trilinear interpolation aligns tensor dimensions prior to skip-connection concatenation. The architecture employs no batch normalization, dropout, or other explicit regularization layers.
The framework obtains the reconstructed channel as
@@
Training uses the Adam optimizer [31] with fixed learning rate $10^{-3}$ and batch size 2 (unless otherwise stated). The procedure uses no learning-rate decay or warm-up. It applies early stopping only to the ResNet baseline, leaving it disabled by default; the hybrid Tensor–NN trains for a fixed number of epochs. The implementation initializes network weights using PyTorch default schemes, corresponding to Kaiming/He uniform initialization for convolutional and transposed convolutional layers, with biases set to zero when present [32].
The training procedure applies no explicit data augmentation. Random channel generation provides training variability, supplemented when enabled by sampling pilot ratios and SNR values over predefined ranges (multi-pilot and multi-SNR training). The hybrid Tensor–NN trains for a fixed number of epochs without early stopping.
@@ 4.4.2 Residual Structure and Architectural Rationale
The U-Net architecture matches this residual structure well: its encoder captures coarse global context through progressive downsampling, while skip connections preserve fine-scale localized information. This design enables efficient reconstruction of multi-scale residual patterns with a small parameter budget, a property proven effective in inverse problems and denoising tasks involving structured residuals.
@@ 5 Simulation Results
This section evaluates pilot-limited wideband MIMO channel estimation under synthetic and ray-tracing-based propagation models through four experiments. Experiment 1 benchmarks Tucker completion against LS, LMMSE, and OMP baselines across pilot densities under a controlled synthetic channel at fixed SNR. Experiment 2 isolates the performance gap between CP and Tucker decompositions as a function of SNR under the synthetic specular channel model. Experiment 3 validates the hybrid Tensor–NN framework on DeepMIMO ray-tracing data, assessing generalization beyond the synthetic model. Experiment 4 characterizes sample complexity by empirically mapping recovery thresholds in the $(L,\rho)$ plane. Across all scenarios, the goal is to quantify how intrinsic tensor structure—in particular, low multilinear rank induced by sparse scattering—enables reliable reconstruction of the wideband channel tensor $\mathcal{H}$ from noisy and incomplete pilot observations.